# Phase transition on the fluctuation of the structure of random walk ranges

### v1

@@ 6.3.1. The large scales
Namely, it is the amount of times that $S_{t}$ lies in the ball $\mathcal{B}_{r}$.
@@
By using the local central limit theorem, one observes that the probability of a point $S_{t}$ being in the ball $\mathcal{B}_{r}$ is of order $\frac{r^{4}}{t^{2}}$. If we sum up this probability over all times $t$ from $n/(\log n)^{4}$ and $n$, we see that the expectation of the local time within this ball is $\frac{r^{4}(\log n)^{4}}{n}$.
@@
$\text{Cap}(\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})\leq|\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n}|$
$\leq|L(\mathcal{B}_{\sqrt{n/K_{i}}})|\leq\frac{n}{(K_{i})^{1+\delta}}.$
@@ 6.3.2. The smallest scales
In [12, Lemma 4.6], we have that these are $c>0$ and $C<\infty$ such that
$\mathbb{E}\left[\exp\left(c\frac{(\log n)^{2}}{n}({\rm Cap}(S[0,n])-\mathbb{E}[{\rm Cap}(S[0,n])])\right)\right]\leq C$
@@
In order to discuss the properties of the parts of the random walk inside each annulus, we need to consider the following stopping times.
Define
$T^{i}_{k}=\inf\{t\geq\tau^{i}_{k-1}:S_{t}\leq\sqrt{n/K_{i}}\},$
$\tau^{i}_{k}=\inf\{t\geq T^{i}_{k}:S_{t}\geq 2\sqrt{n/K_{i}}\}.$
These are a sequence of times indicating when the random walk hits that ball of radius $\sqrt{n/K_{i}}$ and later details when it escapes said ball.
The following lemma will ensure that with high probability that there are not too many times in which the random walk will make an excursion into the ball of radius $\sqrt{n/K_{i}}$. Furthermore, each excursion into this ball will lie very close to its expected time.
@@
Let $M^{i}=\sup\{k:\tau^{i}_{k}\leq n\}$.
Let $\mathcal{H}^{i}$ be the event that
@@
where we recall the notation $L(\mathcal{B}_{r})$ from Lemma 6.10.
By using equation (6.10) and Markov’s inequality with $k$ large enough, we would have that
@@
Now we consider the intersection of events,
$\Omega^{i}=\mathcal{S}^{i}\cap\mathcal{G}^{i}\cap\mathcal{H}^{i}.$
By combining the probability bounds from earlier, we have $\mathbb{P}((\Omega^{i})^{c})\ll(\log n)^{-2}.$
Notice that, on the event $\Omega^{i}$, we can show that,
(6.13) ${\rm Cap}\left(\tilde{S}\left[\frac{n}{(\log n)^{4}},n\right]\cap\mathcal{A}_{i,n}\right)\leq\sum_{k=1}^{M^{i}}{\rm Cap}(\tilde{S}[T^{i}_{k},\tau^{i}_{k}])\leq\sum_{k=1}^{M^{i}}{\rm Cap}(\tilde{S}[T^{i}_{k},T^{i}_{k}+\frac{n}{K_{i}^{1-\kappa}}])$
@@
The second inequality used the fact that $|\tau^{i}_{k}-T^{i}_{k}|\leq\frac{n}{K_{i}^{1-\kappa}}$ from the event $\mathcal{H}^{i}$. The third inequality used the bound on the capacity from the event $\mathcal{G}^{i}$. The fifth inequality used that $|\tau^{i}_{k}-T^{i}_{k}|\geq\frac{1}{K_{i}^{1-\kappa}}$ from the event $\mathcal{H}^{i}$. Note that the sum of these differences is a lower bound on the amount of times in the intersection between $\tilde{S}[\frac{n}{(\log n)^{4}},n]$ and $\mathcal{B}_{2\sqrt{n/K_{i}}}$. The final inequality used the event $\mathcal{S}^{i}$.
@@
By setting $3(2\kappa)+\delta+\chi\leq\epsilon/2$, we see that the line above will be less than $(\log n)^{-\epsilon/2}$. We now consider $\hat{\Omega}:=\bigcap_{i=1}^{11}\Omega^{i}$, we must have $\mathbb{P}(\hat{\Omega}^{c})\ll(\log n)^{-2}$. Furthermore, we must also have
@@
Now, consider the intersection $\tilde{\Omega}\cap\hat{\Omega}$, where we recall $\tilde{\Omega}$ from equation (6.11). We must have that $\mathbb{P}((\tilde{\Omega}\cap\hat{\Omega})^{c})\ll(\log n)^{-2}$ and, in addition, the right-hand side of equation (6.8) can be bounded by $(\log n)^{-c}$ for some $c>0$. This completes the proof of Lemma 6.9. ∎
@@ 6.4. An analysis of $X_{n}-\mathbb{E}[X_{n}]$
In this section, we will prove that $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}\to 0$ in distribution. We will first start by proving upper and lower bounds on $\mathrm{Var}(X_{n})$. Note that an upper bound on $\mathrm{Var}(X_{n})$ was already derived in Lemma 6.6. We now derive the corresponding lower bound for $\mathrm{Var}(X_{n})$. Is this subsection, we will give the estimate only for cut points and graph distance.
@@
Notice, now that since $\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}]$ is a sum of i.i.d. terms where each individual term has variance at most $\frac{n^{2}}{(\log n)^{2-o(1)}}$, we can compute its variance as
@@
Let $X_{n}$ denote the number of cut points, or the graph distance. We have that as $n\to\infty$,
@@
We proceed with the same dyadic decomposition as we have done in the previous proof.
@@
6.5. An analysis of the effective resistance
In this subsection, we present, for independent interest, a similar analysis done in the previous sections when we impose weaker conditions on the bounds we obtain for the variance and the expectation for the cross terms under consideration in $d=4$. Namely, we will assume that the cross-terms are very heavy tailed and that we have some almost optimal upper and lower bounds for the variance up to powers of $\log n$. In particular, we will give the proof of the desired result for the effective resistance in $d=4$ in Subsection 6.5.2. The following assumption will detail what we assume about the cross terms under consideration.
Before this, in Subsection 6.5.1, we will give an estimate for the effective resistance, which assures that Assumption 6.17 parts (2) (3) will hold.
@@ 6.5.1. Upper bound for the cross term of the effective resistance
We will give the upper bound for the second moment of the cross term of the effective resistance $E[E_{n}^{2}]$.
@@
Remark 6.16.
The constant that appears on the right-hand side of equation (6.15) is not the best possible (rather it should be improved to 1o(1)1-o(1)). However, obtaining the optimal constant is not necessary for the estimates that follow.
@@
Recall $\Omega$ in Lemma 6.3. On $\Omega^{c}$, we have that,
@@
Hence, we obtain (6.14). Notice, also that since $E_{n}$ is deterministically less than $n$, we must have that,
@@
By Lemmas 6.1 and 6.15, we obtain the result, which corresponds to (2) (3) for Assumption 6.17 for $b_{1}=2+o(1)$ and $b_{2}=1-o(1)$. ∎
@@ 6.5.2. General arguments
We let $X[a,b]$ for $a,b\in\mathbb{Z}^{+},b>a$ be a collection of random variables satisfying the following properties. First, define $E_{n}:=X_{n}+X[n,2n]-X_{2n}$
@@
    If $a,b]\cap[c,d]=\emptyset$ independent, then $X[a,b]$ is independent of $X[c,d]$.
@@
    There is a set $\Omega$ with probability $O(\frac{\log\log n}{\log n})$ such that
@@
In order to characterize this, we need to first define the associated function $g(n)$ as as follows:
@@
To get the third inequality, we used the fact that $E_{n}^{(k_{2},l_{2})}$ must be positive. Thus, under the condition that $E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$ is negative, the expectation of $\mathbb{E}[(E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}])^{2}\mathbbm{1}[E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}]<0]]^{1/2}\leq\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$. The second to last inequality used the definition of $g$ along with equation (6.16) for $n$ large enough. Finally, the last inequality used the fact that $\frac{n}{(\log n)^{k}}$ is increasing for $n$ large enough as well as the definition of $g(n,K)$.
@@
The factor of $o(1)$ is to compensate for possible changes of the constant factor in front. We remark that due to these computations, there is a gap of of a factor of $(\log n)^{1/2}$ between the positive term above and the negative contribution from equation (6.17). This will ensure a sufficiently good lower bound for the expectation of $\overline{\mathcal{E}}$. Indeed, we have that,
@@
In the first inequality, we used the fact that $\mathcal{I}$ is a sum of independent random variables. Since $g(n,K)$ lies between $b_{2}+o(1)$ and $b_{1}+o(1)$, the final inequality must be true for large enough $n$.
Notice now, that we have, by the Cauchy-Schwarz inequality,
@@
  • If there exists an infinite sequence of $m$ such that we have $f(n2^{-K})$ is greater than $g(n,K)+\frac{1}{4}$, then we have $f(n)\leq g(n,K)-o(1)$. For such $n$, we will be able to show that $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ will go to 0.
@@
Let $X_{n}$ be a series of random variables that satisfies the assumptions of 6.17. We have that there exists a subsequence such that as $n\to\infty$,
@@
Here, $\tilde{X}_{(K,l)}=X[(l-1)n2^{-K},ln2^{-K}]$ is a collection of independent copies with the same distribution as $X_{n2^{-K}}$ and for each given $K$, $E_{n}^{(k,l)}$ is a collection of independent random variables with the same distribution as $E_{n2^{-k}}$.
@@
By the third Assumption 6.17, we see that $\Omega$ occurs with probability 1o(1)1-o(1). We define $\mathcal{E}=-\sum_{k=1}^{K}\sum_{l=1}^{2^{k}}E_{n}^{(k,l)}$ and observe that,
@@
In the last equation, we used equation (6.16) to bound $\mathbb{E}[E_{n}^{(k,l)}]^{2}$ relative to $\mathbb{E}[(E_{n}^{(k,l)})^{2}]$ and we placed irrelevant constant factors into an $o(1)$ exponent of $\log n$ in the denominator.
In addition, we have that,
@@
when $n$ is sufficiently large. Indeed, we used that for $n$ large and fixed $K$, we can bound $\frac{1}{(\log n2^{-K})^{f(n2^{-K})}}\leq\frac{1}{(\log n)^{g(n,K)+\frac{1}{4}+o(1)}}$ where the $o(1)$ term in the exponent is introduced to from $\log n2^{-K}\to\log n$ in the denominator. Recall, we assume $f(n2^{-K})\geq g(n,K)+\frac{1}{4}.$ In the numerator, we bounded $(\log n)^{f(n)}$ by $(\log n)^{g(n,K)+o(1)}$.
Combining this estimate of equation (6.22), we see that
@@
Thus, there will be some set $\Omega^{\prime}$, which will still hold with probability 1o(1)1-o(1) such that,
@@
Furthermore, recall from equation (6.21) that the sums of the cross terms $\mathcal{E}$ satisfies,
@@
Again, there will be a further subset $\Omega^{\prime\prime}\subset\Omega^{\prime}$ that will still hold with probability 1o(1)1-o(1) on which we have additionally that $|\mathcal{E}-\mathbb{E}[\mathcal{E}]|<\epsilon\sqrt{\mathrm{Var}(X_{n})}$. Thus, on $\Omega^{\prime\prime}$, we have $|X_{n}-\mathbb{E}[X_{n}]|<2\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
@@
Case 2a: There does not exists an infinite sequence of $n$ such that $f(n)-f(n2^{-K})\leq\frac{1}{\log\log n}$
@@
Now, since $\sum_{m=1}^{\infty}\frac{1}{\log\log(a2^{mK})}$ diverges, this must mean that $f(a2^{mK})$ is unbounded from above. This contradicts our assumption that $f(n)\leq g(n,K)+\frac{1}{4}$ for $x$ large enough. Thus, we do not have to worry about this case.
@@
$\leq$ $(\log n)^{\frac{1}{\log\log n}}\left(\frac{\log n}{\log n-K\log 2}\right)^{2+\frac{1}{4}}\leq 2e,$
where the final inequality will hold for large enough $m$. Notice that we also used that $f(n2^{-K})\leq g(n,K)+\frac{1}{4}\leq\frac{3}{2}+\frac{1}{4}+o(1)$, by our assumption that eventually all $f(n)\leq g(n,K)+\frac{1}{4}$ and our upper bound on $g(n,K)$.
@@
Again, here we used that first we are considering an $m$ such that first we have $f(n)\leq f(n2^{-K})+\frac{1}{\log\log n}$. Secondly, under the general assumption of case 2, we have $f(n2^{-K})+\frac{1}{\log\log n}\leq g(n,K)+\frac{1}{4}+\frac{1}{\log\log n}\leq g(n,K)+1+o(1)$ for $n$ large enough. Thus, we have $f(n)\leq g(n,K)+1+o(1)$ for some infinite sequence of $n$. Thus (again by Markov’s inequality on the set $\Omega$), with probability $o(1)$, we have that $|\mathcal{E}-\mathbb{E}[\mathcal{E}]|\geq\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
In total, we see that with probability at most $o(1)+2e2^{-K}\epsilon^{-2}$, we have that
@@
$|X_{n_{i}}-\mathbb{E}[X_{n_{i}}]|\leq\epsilon\sqrt{\mathrm{Var}(X_{n_{i}})}$
@@
  • [6] Y. Chang. Two observations on the capacity of the range of simple random walk on Z3 and Z4. Electron. Com. Probab. 22 (2017), No. 25, 9 pp.

### v2

@@ 6.3.1. The large scales
Namely, it is the number of times that $S_{t}$ lies in the ball $\mathcal{B}_{r}$.
@@
By using the local central limit theorem, the probability that a point, $S_{t}$, lies in the ball $\mathcal{B}_{r}$ is of order $\frac{r^{4}}{t^{2}}$. If we sum up this probability over all times $t$ from $n/(\log n)^{4}$ and $n$, we see that the expectation of the local time within this ball is $\frac{r^{4}(\log n)^{4}}{n}$.
@@
$\text{Cap}(\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n})\leq|\tilde{S}[\frac{n}{(\log n)^{4}},n]\cap\mathcal{A}_{i,n}|\leq|L(\mathcal{B}_{\sqrt{n/K_{i}}})|\leq\frac{n}{(K_{i})^{1+\delta}}.$
@@ 6.3.2. The smallest scales
In [12, Lemma 4.6], we have that there is $c>0$ such that
$\limsup_{n\to\infty}\mathbb{E}\left[\exp\left(c\frac{(\log n)^{2}}{n}({\rm Cap}(S[0,n])-\mathbb{E}[{\rm Cap}(S[0,n])])\right)\right]<\infty$
@@
In order to discuss the properties of the parts of the random walk inside each annulus, we need to consider the following stopping times. Define
$T^{i}_{k}=\inf\{t\geq\tau^{i}_{k-1}:S_{t}\leq\sqrt{n/K_{i}}\},\quad\tau^{i}_{k}=\inf\{t\geq T^{i}_{k}:S_{t}\geq 2\sqrt{n/K_{i}}\}.$
These are a sequence of times that indicate when the random walk hits that ball of radius $\sqrt{n/K_{i}}$, and later details when it escapes that ball.
The following lemma will ensure that, with high probability, there are not too many times in which the random walk will make an excursion into the ball of radius $\sqrt{n/K_{i}}$. Furthermore, each excursion into this ball will lie very close to its expected time.
@@
Let $M^{i}=\sup\{k:\tau^{i}_{k}\leq n\}$. Let $\mathcal{H}^{i}$ be the event that
@@
where we recall the notation $L(\mathcal{B}_{r})$ from Lemma 6.10. By using equation (6.10) and Markov’s inequality with $k$ large enough, we would have that
@@
Now, we consider the intersection of events $\Omega^{i}:=\mathcal{S}^{i}\cap\mathcal{G}^{i}\cap\mathcal{H}^{i}$. By combining the probability bounds from earlier, we have $\mathbb{P}((\Omega^{i})^{c})\ll(\log n)^{-2}$. Notice that, on the event $\Omega^{i}$, we can show that,
${\rm Cap}\left(\tilde{S}\left[\frac{n}{(\log n)^{4}},n\right]\cap\mathcal{A}_{i,n}\right)\leq\sum_{k=1}^{M^{i}}{\rm Cap}(\tilde{S}[T^{i}_{k},\tau^{i}_{k}])\leq\sum_{k=1}^{M^{i}}{\rm Cap}(\tilde{S}[T^{i}_{k},T^{i}_{k}+\frac{n}{K_{i}^{1-\kappa}}])$
@@
The second inequality used the fact that $|\tau^{i}_{k}-T^{i}_{k}|\leq\frac{n}{K_{i}^{1-\kappa}}$ from the event $\mathcal{H}^{i}$. The third inequality used the bound on the capacity from the event $\mathcal{G}^{i}$. The fifth inequality used that $|\tau^{i}_{k}-T^{i}_{k}|\geq\frac{1}{K_{i}^{1-\kappa}}$ from the event $\mathcal{H}^{i}$. Note that the sum of these differences is a lower bound on the number of times in the intersection between $\tilde{S}[\frac{n}{(\log n)^{4}},n]$ and $\mathcal{B}_{2\sqrt{n/K_{i}}}$. The final inequality used the event $\mathcal{S}^{i}$.
@@
By setting $11(2\kappa+\delta)+\chi\leq\epsilon/2$, we see that the line above will be less than $(\log n)^{-\epsilon/2}$. We now consider $\hat{\Omega}:=\bigcap_{i=1}^{11}\Omega^{i}$, we must have $\mathbb{P}(\hat{\Omega}^{c})\ll(\log n)^{-2}$. Furthermore, we must also have
@@
Now, consider the intersection $\tilde{\Omega}\cap\hat{\Omega}$, where we recall $\tilde{\Omega}$ from equation (6.11). We must have that $\mathbb{P}((\tilde{\Omega}\cap\hat{\Omega})^{c})\ll(\log n)^{-2}$ and, in addition, that the right-hand side of equation (6.8) can be bounded by $(\log n)^{-c}$ for some $c>0$. This completes the proof of Lemma 6.9. ∎
@@ 6.4. An analysis of $X_{n}-\mathbb{E}[X_{n}]$
In this section, we will prove that $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}\to 0$ in distribution. We will first start by proving upper and lower bounds on $\mathrm{Var}(X_{n})$. Note that an upper bound on $\mathrm{Var}(X_{n})$ was already derived in Lemma 6.6. We now derive the corresponding lower bound for $\mathrm{Var}(X_{n})$.
@@
Notice, that since $\sum_{j=0}^{(\log n)^{\epsilon}-1}X[jn/(\log n)^{\epsilon},(j+1)n/(\log n)^{\epsilon}]$ is a sum of i.i.d. terms where each individual term has variance at most $\frac{n^{2}}{(\log n)^{2-o(1)}}$, we can compute its variance as
@@
Let $X_{n}$ denote the number of cut points or the graph distance. We have that, as $n\to\infty$,
@@
We proceed with the same dyadic decomposition as in the previous proof.
@@
6.5. An Analysis of Variables with heavy tailed Cross Terms
In this subsection, we present, for independent interest, a similar analysis done in the previous sections when we impose weaker conditions on the bounds we obtain for the variance and the expectation for the cross terms under consideration in $d=4$. Namely, we will assume that the cross terms are very heavy tailed and that we have some almost optimal upper and lower bounds for the variance up to powers of $\log n$. In particular, we will show that the effective resistance in $d=4$ satisfies these assumptions in Subsection 6.5.2. Assumption 6.16 will detail what we assume about the cross terms under consideration.
Before this, in Subsection 6.5.1, we will briefly derive estimates for the effective resistance, which ensure that parts (2) and (3) of Assumption 6.16 will hold.
@@ 6.5.1. Upper bound for the cross term of the effective resistance
We will start by giving the upper bound for the second moment of the cross term of the effective resistance, $E[E_{n}^{2}]$.
@@
Recall $\Omega$ from Lemma 6.3. The same proof will allow us to say that on $\Omega^{c}$, we have that,
@@
Hence, we obtain (6.13). Notice also that since $E_{n}$ is deterministically less than $n$, we must have that,
@@
By Lemmas 6.1 and 6.15, which also hold for the effective resistance, we obtain results which correspond to parts (2) and (3) of Assumption 6.16 for $b_{1}=2+o(1)$ and $b_{2}=1-o(1)$. ∎
@@ 6.5.2. General arguments
We let $X[a,b]$ for $a,b\in\mathbb{Z}^{+},b>a$ be a collection of random variables satisfying the following properties. First, define $E_{n}:=X_{n}+X[n,2n]-X_{2n}.$
@@
    If $a,b]\cap[c,d]=\emptyset$, then $X[a,b]$ is independent of $X[c,d]$.
@@
    There is an event $\Omega$ with probability $O(\frac{\log\log n}{\log n})$ such that
@@
In order to characterize this, we need to first define the associated function $g(n)$ as follows:
@@
To get the third inequality, we used the fact that $E_{n}^{(k_{2},l_{2})}$ must be positive. Thus, under the condition that $E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$ is negative, the expectation of $\mathbb{E}[(E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}])^{2}\mathbbm{1}[E_{n}^{(k_{2},l_{2})}-\mathbb{E}[E_{n}^{(k_{2},l_{2})}]<0]]^{1/2}\leq\mathbb{E}[E_{n}^{(k_{2},l_{2})}]$. The second-to-last inequality used the definition of $g$ along with the equation (6.15) for $n$ large enough. Finally, the last inequality used the fact that $\frac{n}{(\log n)^{k}}$ is increasing for $n$ large enough, as well as the definition of $g(n,K)$.
@@
The factor of $o(1)$ is to compensate for possible changes of the constant factor in front. We remark that due to these computations, there is a gap of a factor of $(\log n)^{1/2}$ between the positive term above and the negative contribution from equation (6.16). This will ensure a sufficiently good lower bound for the expectation of $\overline{\mathcal{E}}$. Indeed, we have that,
@@
In the first inequality, we used the fact that $\mathcal{I}$ is a sum of independent random variables. Since $g(n,K)$ lies between $b_{2}+o(1)$ and $b_{1}+o(1)$, the final inequality must be true for a large enough $n$.
Finally, notice that, by the Cauchy-Schwarz inequality,
@@
  • If there exists an infinite sequence of $m$ such that $f(n2^{-K})$ is greater than $g(n,K)+\frac{1}{4}$, then $f(n)\leq g(n,K)-o(1)$. For such $n$, we will be able to show that $\frac{X_{n}-\mathbb{E}[X_{n}]}{\sqrt{\mathrm{Var}(X_{n})}}$ will go to 0.
@@
Let $X_{n}$ be a series of random variables that satisfy the assumptions of 6.16. We have that there exists a subsequence such that as $n\to\infty$,
@@
Here, $\tilde{X}_{(K,l)}=X[(l-1)n2^{-K},ln2^{-K}]$ is a collection of independent copies with the same distribution as $X_{n2^{-K}}$, and for each given $K$, $E_{n}^{(k,l)}$ is a collection of independent random variables with the same distribution as $E_{n2^{-k}}$.
@@
By the third part of Assumption 6.16, we see that $\Omega$ occurs with probability 1o(1)1-o(1). We define $\mathcal{E}=-\sum_{k=1}^{K}\sum_{l=1}^{2^{k}}E_{n}^{(k,l)}$ and observe that,
@@
In the last equation, we used equation (6.15) to bound $\mathbb{E}[E_{n}^{(k,l)}]^{2}$ relative to $\mathbb{E}[(E_{n}^{(k,l)})^{2}]$, and we placed irrelevant constant factors into an $o(1)$ exponent of $\log n$ in the denominator. In addition, we have that,
@@
when $n$ is sufficiently large. Indeed, we used that for $n$ large and fixed $K$, we can bound $\frac{1}{(\log n2^{-K})^{f(n2^{-K})}}\leq\frac{1}{(\log n)^{g(n,K)+\frac{1}{4}+o(1)}}$, where the $o(1)$ term in the exponent comes from the change $\log n2^{-K}\to\log n$ in the denominator. Recall, we assume $f(n2^{-K})\geq g(n,K)+\frac{1}{4}.$ In the numerator, we bounded $(\log n)^{f(n)}$ by $(\log n)^{g(n,K)+o(1)}$.
Combining this estimate with equation (6.21), we see that
@@
Thus, there will be some set $\Omega^{\prime}$, which will still hold with probability 1o(1)1-o(1), such that,
@@
Furthermore, recall from equation (6.20) that the sums of the cross terms $\mathcal{E}$ satisfy,
@@
Again, there will be a further sub-event $\Omega^{\prime\prime}\subset\Omega^{\prime}$, that will still hold with probability 1o(1)1-o(1), on which we have that $|\mathcal{E}-\mathbb{E}[\mathcal{E}]|<\epsilon\sqrt{\mathrm{Var}(X_{n})}$. Thus, on $\Omega^{\prime\prime}$, we have $|X_{n}-\mathbb{E}[X_{n}]|<2\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
@@
Case 2a: There does not exist an infinite sequence of $n$ such that $f(n)-f(n2^{-K})\leq\frac{1}{\log\log n}$
@@
Now, since $\sum_{m=1}^{\infty}\frac{1}{\log\log(a2^{mK})}$ diverges, this must mean that $f(a2^{mK})$ is unbounded from above. This contradicts our assumption that $f(n)\leq g(n,K)+\frac{1}{4}$ for $n$ large enough. Thus, we do not have to worry about this case.
@@
$\leq$ $(\log n)^{\frac{1}{\log\log n}}\left(\frac{\log n}{\log n-K\log 2}\right)^{b_{2}+\frac{1}{4}}\leq 2e,$
where the final inequality will hold for large enough $n$. Notice that we also used $f(n2^{-K})\leq g(n,K)+\frac{1}{4}\leq b_{2}+\frac{1}{4}+o(1)$, by our assumption that eventually all $f(n)\leq g(n,K)+\frac{1}{4}$ and our upper bound on $g(n,K)$.
@@
Again, here we used that first we are considering an $n$ such that first we have $f(n)\leq f(n2^{-K})+\frac{1}{\log\log n}$. Secondly, under the general assumption of case 2, we have $f(n2^{-K})+\frac{1}{\log\log n}\leq g(n,K)+\frac{1}{4}+\frac{1}{\log\log n}\leq g(n,K)+1+o(1)$ for $n$ large enough. Thus, we have $f(n)\leq g(n,K)+1+o(1)$ for some infinite sequence of $n$. Thus (again by Markov’s inequality on the event $\Omega$), with probability $o(1)$, we have $|\mathcal{E}-\mathbb{E}[\mathcal{E}]|\geq\epsilon\sqrt{\mathrm{Var}(X_{n})}$.
In total, we see that with probability at most $o(1)+2e2^{-K}\epsilon^{-2}$,
@@
$|X_{n_{i}}-\mathbb{E}[X_{n_{i}}]|\leq\epsilon\sqrt{\mathrm{Var}(X_{n_{i}})}.$
@@
  • [6] Y. Chang. Two observations on the capacity of the range of simple random walk on $\mathbb{Z}^{3}$ and $\mathbb{Z}^{4}$. Electron. Com. Probab. 22, No. 25, 2017.